# Singularities of solutions of nonlocal nonlinear equations

### v1

@@ 1. Introduction
The classical results due to Serrin [33, 34] deal with the local behavior of weak solutions of quasilinear equations in divergence form, whose model equation is the $p$-Laplace equation $-\Delta_{p}u=0$ for $p>1$. Regarding the regular behavior of solutions, he established the local boundedness, the weak Harnack inequality, the Harnack inequality and Hölder continuity by adapting Moser’s iteration technique [29, 30] to the quasilinear setting. These regularity results were then employed to investigate the singularity of solutions, leading to the removable singularity theorem and the isolated singularity theorem. He first showed that a solution $u$ in $\Omega\setminus E$ can be extended as a solution of the same equation in the whole domain $\Omega$, provided that $E$ is negligible in the sense of capacity and that $u$ satisfies certain integrability conditions. Furthermore, he proved that singular solutions near an isolated singularity must behave like the fundamental solution of the $p$-Laplace equation.
The goal of this paper is to establish a nonlocal counterpart of the results in the celebrated papers [33, 34]. In particular, we are interested in the nature of removable singularities of solutions and the behavior of a solution near an isolated singularity in the nonlocal nonlinear framework. It is noteworthy that the local regularity of weak solutions of nonlocal nonlinear equations has been relatively well studied in various works, including those by Di Castro–Kuusi–Palatucci [17, 18] and Cozzi [15].
Let $n\in\mathbb{N}$, $0<s<1<p<\infty$, $0<\varepsilon<sp$ and $\Lambda\geq 1$, and assume that $sp\leq n$. Let $\Omega$ be a nonempty open set in $\mathbb{R}^{n}$. We are concerned with weak solutions $u$ of the nonlocal nonlinear equation
@@
with nonnegative functions $b_{1},b_{2}\in L^{n/(sp-\varepsilon)}(\Omega)$. The principal model equation is the fractional $p$-Laplace equation $(-\Delta_{p})^{s}u=0$.
Our first main theorem on the removability singularity reads as follows. See Section 2 for the definitions of weak solutions and sets of capacity zero.
@@
A few remarks on Theorem 1.1 are in order. In contrast to the local case, $u$ is a priori defined on the whole of $\mathbb{R}^{n}$ in Theorem 1.1. Moreover, since $E$ is the set of measure zero by Lemma 2.15, the condition $u\in L^{(1+\varepsilon)\theta}(\Omega\setminus E)$ in Theorem 1.1 is equivalent to $u\in L^{(1+\varepsilon)\theta}(\Omega)$. When $q=p$, this condition is understood as $u\in L^{\infty}(\Omega)$. Note however that the requirement $u\in W^{s,p}_{\mathrm{loc}}(\Omega\setminus E)\cap L^{p-1}_{sp}(\mathbb{R}^{n})$ for $u$ to be a weak solution cannot be replaced by $u\in W^{s,p}_{\mathrm{loc}}(\Omega)\cap L^{p-1}_{sp}(\mathbb{R}^{n})$ (see Definition 2.5). This makes the problem significantly more challenging.
If $E$ is a singleton, then $E$ is of $(s,p)$-capacity zero by Lemma 2.16. Thus, Theorem 1.1 implies the following removable isolated singularity theorem under a suitable growth condition in the subcritical case.
@@
Our second main theorem is concerned with the classification of the behavior of solutions near an isolated singularity. We begin with the subcritical case. Here and below, $u\simeq v$ means that $C_{1}\leq u/v\leq C_{2}$ for some positive constants $C_{1}$ and $C_{2}$. ($u\lesssim v$ and $u\gtrsim v$ can be defined in same way.)
@@
Let us briefly summarize the historical background. Singularities of solutions of local equations have been of great interest; just to name a few, we refer to [6, 10, 21] for linear equations, [31, 33, 34] for quasilinear equations and [1, 25, 26] for fully nonlinear equations. All results on isolated singularities share the essence that if $u$ is a solution in the punctured ball, then either $u$ has a removable singularity or $u$ behaves like a fundamental solution of the associated local operator. We also refer to [9, 11, 24] for scalar curvature equations and [2, 3, 4, 20, 28] for other semilinear equations, where the asymptotic behavior is captured by different functions than fundamental solutions. See also comprehensive monographs [32, 35] and references therein.
Nevertheless, the body of literature on singularities of solutions of nonlocal equations is relatively smaller than that of local ones. Caffarelli–Jin–Sire–Xiong [7], Chen–Quaas [12], Chen–Véron [13], Chen–Li–Ou [14] and Yang–Zou [36, 37] investigated the local behavior of nonnegative solutions of fractional semilinear equations given by
{$(-\Delta)^{s}u&=u^{\alpha}&&\text{in $B_{1}\setminus\{0\}$},\\ u&=0&&\text{in $\mathbb{R}^{n}\setminus B_{1}$},\end{aligned}\right.$
for different ranges of $\alpha$. We point out that these papers rely heavily on the Caffarelli–Silvestre extension [8] that guarantees the validity of various tools in the local setting. In fact, the condition that $u$ vanishes on the complement of $B_{1}$ removes the effect of the nonlocal tail term, which encodes the long-range interaction of $u$. Moreover, Li–Liu–Wu–Xu [27] established the Bôcher-type theorems for nonnegative solutions of fractional linear equation, which characterizes isolated singularities by Dirac mass $\delta_{0}$ in the distribution sense.
@@
Let us finally illustrate the strategies to deduce the principal results of the present paper. In contrast to the simple display of our main theorems, their proofs are quite intricate and rely heavily on advanced tools from nonlocal regularity theory. As mentioned earlier, the major difficulty in proving the removability theorem (Theorem 1.1) arises from the lack of the regularity of $u$ in the whole domain $\Omega$. To address this difficulty, we develop the De Giorgi–Nash–Moser theory tailored to our situation. Indeed, in Section 3, we obtain Moser-type Caccioppoli estimates that contain a cut-off function $\bar{\eta}$ vanishing in a neighborhood of $E$, which is introduced to avoid the possible singularity on $E$, and some powers of weak (sub-)solutions $u$. On the one hand, we need a polynomial of $u$ with a certain growth rate to use the integrability condition given in Theorem 1.1. On the other hand, we also need a polynomial of $u$ with an arbitrarily large power in order for Moser’s iteration to work. By taking a test function as the minimum of these two polynomials, we arrive at the desired Caccioppoli estimate.
@@
The proof of the isolated singularity theorem (Theorem 1.3 or Theorem 1.5) is carried out by combining several steps as follows. If a weak solution $u$ has a non-removable isolated singularity at 00, then
@@
The paper is organized as follows. In Section 2, we summarize the notations and preliminary results regarding functional inequalities, weak solutions and capacities. Section 3 consists of local regularity estimates, namely, various Caccioppoli-type estimates, Hölder regularity and the Harnack inequality. The proofs for Theorem 1.1 and Theorem 1.3 are presented in Section 4 and Section 5, respectively.
@@ 2.1. Functional inequalities
Some functional inequalities, such as the fractional Sobolev inequality and Hölder’s inequality, are presented here. We begin with the fractional Sobolev inequalities. See Di Nezza–Palatucci–Valdinoci [19, Theorems 6.5 and 6.7] for their proofs.
@@
(Cozzi [15, Lemma 4.6]) Let $\tilde{s}\in(0,s)$, $\tilde{p}\in(1,p)$ and $B=B_{r}(x_{0})$. Then there exists a constant $C=C(n,s,\tilde{s},p,\tilde{p})>0$ such that
@@ 2.2. Weak solution
where $\Lambda$ is the ellipticity constant given as in (1.2).
@@ 2.3. $(s,p)$-capacity
As pointed out in Björn–Björn–Kim [5, Section 5], (2.7) and (2.8) do not change the definition of capacities for compact sets.
The following estimates for the $(s,p)$-capacity of balls are known in Kim–Lee–Lee [22, Lemma 2.17]. By $A\simeq B$ it means that $B/C\leq A\leq A/C$ for some constant $C\geq 1$ independent of the quantities $A$ and $B$.
@@
$\mathrm{cap}_{s,p}(\overline{B_{r}},B_{R})\simeq\begin{cases}\left(r^{-\frac{n-sp}{p-1}}-R^{-\frac{n-sp}{p-1}}\right)^{1-p}&\text{if }sp<n,\\ (\log(R/r))^{1-p}&\text{if }sp=n,\end{cases}$
@@
By using Lemma 2.12, we can also obtain the estimates for the $(s,p)$-capacity of the annulus as follows.
Lemma 2.13.
Let $B_{2r}\Subset\Omega$. Then
$\mathrm{cap}_{s,p}(\overline{B_{r}}\setminus B_{r/2},\Omega)\simeq\mathrm{cap}_{s,p}(\overline{B_{r}},\Omega)\simeq\begin{cases}r^{n-sp}&\text{if}~{}sp<n,\\ (-\log r)^{1-p}&\text{if}~{}sp=n,\end{cases}$
where the comparable constants do not depend on $r$.
@@
Without loss of generality, we may assume that $\Omega=B_{R}$ for some $R>0$. In view of Lemma 2.12 and the monotonicity of capacity, it suffices to prove that
@@
Let $\Omega$ be a bounded open set and let $\varepsilon>0$. Since $\mathrm{cap}_{s,p}(E\cap\Omega,\Omega)=0$, there exists an open neighborhood $G\subset\Omega$ of $E\cap\Omega$ such that $\mathrm{cap}_{s,p}(G,\Omega)<\varepsilon$. Let $K\subset G$ be a compact set and take $\varphi\in W(K,\Omega)$ such that $[\varphi]_{W^{s,p}(\mathbb{R}^{n})}^{p}\leq\mathrm{cap}_{s,p}(K,\Omega)+\varepsilon<2\varepsilon$. Thus, the fractional Poincaré inequality (see e.g. Björn–Björn–Kim [5, (5.6)]) shows that
@@
It follows from [5, Lemma 5.5] that $sp\leq n$ if and only if $\mathrm{cap}_{s,p}(\{x_{0}\},\Omega)=0$ for every bounded open $\Omega$ containing $x_{0}$. Since $\mathrm{cap}_{s,p}(\{x_{0}\},\cdot)$ is decreasing with respect to the set inclusion, the desired result follows. ∎
We close the section with the following lemma, which is useful when the $(s,p)$-capacity of an arbitrary set is approximated by the seminorms of functions.
Lemma 2.17.
Let $E\subset\mathbb{R}^{n}$. Then
(2.9) $\mathrm{cap}_{s,p}(E,\mathbb{R}^{n})=\inf_{u}{[u]_{W^{s,p}(\mathbb{R}^{n})}^{p}},$
where the infimum is taken over all $u\in W^{s,p}(\mathbb{R}^{n})$ such that $u=1$ in a neighborhood of $E$ and $0\leq u\leq 1$ everywhere.
Proof.
The proof goes by the exact same arguments as in that of [5, Lemma 5.7], with the only difference that we now use $[\cdot]_{W^{s,p}(\mathbb{R}^{n})}$ as a norm on $W^{s,p}(\mathbb{R}^{n})$, which is possible by means of the fractional Poincaré inequality. The equality (2.9) for general set $E$ follows immediately in view of (2.8). ∎
@@ 3.1. Caccioppoli-type estimates
In order to state the Caccioppoli-type estimates, we fix $\beta\geq\beta_{0}>0$, $\gamma=\beta+p-1$ and $\gamma_{0}=\beta_{0}+p-1$. For $l>d>0$, we define functions $F,G:[d,\infty)\to[0,\infty)$ by
@@
where the constants $c,C>0$ depending only on $p$.
@@
We may assume that $a>b$. Note that
$L$ $\coloneqq\frac{\gamma}{\beta_{0}}(a-b)^{p-1}(G(A)\eta_{1}^{p}-G(B)\eta_{2}^{p})\geq\frac{\gamma}{\beta_{0}}(A-B)^{p-1}(G(A)\eta_{1}^{p}-G(B)\eta_{2}^{p})$
$=\frac{\gamma}{\beta_{0}}(A-B)^{p-1}(G(A)-G(B))\eta_{2}^{p}+\frac{\gamma}{\beta_{0}}(A-B)^{p-1}G(A)(\eta_{1}^{p}-\eta_{2}^{p})$
$=\frac{\gamma}{\beta_{0}}(A-B)^{p-1}(G(A)-G(B))\eta_{1}^{p}+\frac{\gamma}{\beta_{0}}(A-B)^{p-1}G(B)(\eta_{1}^{p}-\eta_{2}^{p}).$
@@
Let us write $B_{r}=B_{r}(x_{0})$ for simplicity. We define $\varphi=G(\bar{u})(\eta\bar{\eta})^{p}\geq 0$. By Proposition 2.6, $\varphi$ can be used as a test function for (1.1) in $B_{R}\setminus E$. Moreover, since $\mathcal{E}(\bar{u},\varphi)\leq\mathcal{E}(u,\varphi)$ (c.f. the proof of Theorem 1.4 in Di Castro–Kuusi–Palatucci [18]), we have
$\mathcal{E}(\bar{u},\varphi)+\int_{B_{R}\setminus E}b(x,u(x))\varphi(x)\,\mathrm{d}x\leq 0.$
@@
Combining all the estimates above and using Young’s inequality yield
@@
Combining the estimates above yields (3.6) with the exponent $\varepsilon/(sp)$ replaced by $\varepsilon/(2\tilde{s}p)$. The desired estimate (3.3) then follows from [33, Lemma 2] and that $\gamma^{2\tilde{s}p(p-1)/\varepsilon}\leq\gamma^{p-1}+\gamma^{2sp(p-1)/\varepsilon}$ as in the previous case. ∎
@@
Note that Lemma 3.1 and Lemma 3.3 are Moser-type Caccioppoli estimates. We also need the following De Giorgi-type Caccioppoli estimate.
@@
We basically follow the lines of proof of Proposition 8.5 in Cozzi [15] to estimate the left-hand side of (3.9), but we keep the cut-off functions $\eta$ and $\bar{\eta}$. Indeed, a slight modification of [15, (8.4)–(8.6) and (8.10)–(8.12)] shows that
@@
The rest of the proof goes in the same way as before. ∎
@@
When $E=\emptyset$ in Lemma 3.4, we obtain the following standard Caccioppoli estimate. It shows that any weak subsolution of (1.1) belongs to the so-called (fractional) De Giorgi class, cf. Section 6.1 in Cozzi [15]. Note that such Caccioppoli estimates were obtained for equations similar to (1.1) in [15, Proposition 8.5], but the equation (1.1) was not covered.
@@
The proof goes as in those of Lemma 3.4 and Lemma 3.5, with the only difference that we take $\bar{\eta}\equiv 1$. ∎

### v2

@@ 1. Introduction
The classical results due to Serrin [39, 40] deal with the local behavior of weak solutions of quasilinear equations in divergence form, whose model equation is the $p$-Laplace equation $-\Delta_{p}u=0$ for $p>1$. Regarding the regular behavior of solutions, he established the local boundedness, the weak Harnack inequality, the Harnack inequality and Hölder continuity by adapting Moser’s iteration technique [35, 36] to the quasilinear setting. These regularity results were then employed to investigate the singularity of solutions, leading to the removable singularity theorem and the isolated singularity theorem. He first showed that a solution $u$ in $\Omega\setminus E$ can be extended as a solution of the same equation in the entire domain $\Omega$, provided that $E$ is negligible in the sense of capacity and that $u$ satisfies certain integrability conditions. Furthermore, he proved that singular solutions near an isolated singularity must behave like the fundamental solution of the $p$-Laplace equation.
The goal of this paper is to establish a nonlocal counterpart of the results in celebrated papers [39, 40]. In particular, we are interested in the nature of removable singularities of solutions and the behavior of a solution near an isolated singularity in the nonlocal nonlinear framework. It is noteworthy that the local regularity of weak solutions of nonlocal equations has been relatively well studied in various works, including those by Caffarelli–Chan–Vasseur [8] and Kassmann [25] for linear equations and by Di Castro–Kuusi–Palatucci [19, 20] and Cozzi [17] for nonlinear equations.
Let $n\in\mathbb{N}$, $0<s<1<p<\infty$ and $\Lambda\geq 1$, and assume that $sp\leq n$. Let $\Omega$ be a nonempty open set in $\mathbb{R}^{n}$, which may be unbounded. We are concerned with weak solutions $u$ of the nonlocal nonlinear equation
@@
with nonnegative functions $b_{1},b_{2}\in L^{n/(sp-\varepsilon)}(\Omega)$ ($0<\varepsilon<sp$). The principal model equation is the fractional $p$-Laplace equation $(-\Delta_{p})^{s}u=0$.
Our first main theorem on the removable singularity reads as follows. See Section 2 for the definitions of weak solutions and sets of capacity zero.
@@
A few remarks on Theorem 1.1 are in order. In contrast to the local case, $u$ is a priori defined on the whole of $\mathbb{R}^{n}$ in Theorem 1.1. Moreover, since $E$ is the set of measure zero by Lemma 2.14, the condition $u\in L^{(1+\varepsilon)\theta}(\Omega\setminus E)$ in Theorem 1.1 is equivalent to $u\in L^{(1+\varepsilon)\theta}(\Omega)$. When $q=p$, this condition is understood as $u\in L^{\infty}(\Omega)$. Note, however, that the requirement $u\in W^{s,p}_{\mathrm{loc}}(\Omega\setminus E)\cap L^{p-1}_{sp}(\mathbb{R}^{n})$ for $u$ to be a weak solution cannot be replaced by $u\in W^{s,p}_{\mathrm{loc}}(\Omega)\cap L^{p-1}_{sp}(\mathbb{R}^{n})$ (see Definition 2.5). This makes the problem significantly more challenging.
If $E$ is a singleton, then $E$ is of $(s,p)$-capacity zero according to Lemma 2.15. Thus, Theorem 1.1 implies the following removable isolated singularity theorem under a suitable growth condition in the subcritical case.
@@
Our second main theorem is concerned with the classification of the behavior of solutions near an isolated singularity. We begin with the subcritical case. Here and below, $u\eqsim v$ means that $C_{1}\leq u/v\leq C_{2}$ for some positive constants $C_{1}$ and $C_{2}$. ($u\lesssim v$ and $u\gtrsim v$ can be defined in the same way.)
@@
Let us briefly summarize the historical background. Removable singularity and isolated singularity of solutions of local equations have been of great interest; just to name a few, we refer to [7, 12, 23] for linear equations, [37, 39, 40] for quasilinear equations and [2, 30, 31] for fully nonlinear equations. All results on isolated singularities share the essence that if $u$ is a solution in the punctured ball, then either $u$ has a removable singularity or $u$ behaves like a fundamental solution of the associated local operator. We also refer to [11, 13, 29] for the scalar curvature equations and [3, 4, 5, 22, 34] for other semilinear equations, where the asymptotic behavior is captured by functions different from fundamental solutions. See also comprehensive monographs [38, 41] and references therein.
There has been a growing interest in the study on singularities of solutions of nonlocal equations. Li–Wu–Xu [33], Li–Liu–Wu–Xu [32] and Klimsiak [28] established the Bôcher-type theorems for nonnegative solutions of fractional linear equations, which characterize isolated singularities by Dirac mass $\delta_{0}$ in the distribution sense. Moreover, there have been several works focusing on fractional semilinear equations of the form
(1.4) $(-\Delta)^{s}u=u^{\alpha}$
for some $\alpha>1$. Caffarelli–Jin–Sire–Xiong [9], Chen–Quaas [14], Chen–Véron [15], Chen–Li–Ou [16] and Yang–Zou [42, 43] investigated the local behavior of solutions of (1.4) in $B_{1}\setminus\{0\}$ with zero complement data for different ranges of $\alpha$. In particular, for the critical case $\alpha=(n+2s)/(n-2s)$ that corresponds to the fractional Yamabe equation, Jin–de Queiroz–Sire–Xiong [24] described local behavior of solutions with a singular set of fractional capacity zero. Let us also mention the work by Ao–González–Hyder–Wei [1], where a removability result for the equation (1.4) was established.
We point out that most of these papers rely heavily on the Caffarelli–Silvestre extension [10] that guarantees the validity of various tools in the local setting. In fact, the condition that $u$ vanishes on the complement of $B_{1}$ removes the effect of the nonlocal tail term, which encodes the long-range interaction of $u$.
@@
Let us finally illustrate the strategies to deduce the principal results of the present paper. In contrast to the simple display of our main theorems, their proofs are quite intricate and rely heavily on advanced tools from nonlocal regularity theory. As mentioned earlier, the main difficulty in proving the removability theorem (Theorem 1.1) arises from the lack of the regularity of $u$ in the whole domain $\Omega$. To address this difficulty, we develop the De Giorgi–Nash–Moser theory tailored to our situation. Indeed, in Section 3, we obtain Moser-type Caccioppoli estimates that contain a cut-off function $\bar{\eta}$ vanishing in a neighborhood of $E$, which is introduced to avoid the possible singularity on $E$, and some powers of weak (sub-)solutions $u$. On the one hand, we need a polynomial of $u$ with a certain growth rate to use the integrability condition given in Theorem 1.1. On the other hand, we also need a polynomial of $u$ with an arbitrarily large power in order for Moser’s iteration to work. By taking a test function as the minimum of these two polynomials, we arrive at the desired Caccioppoli estimate.
@@
The proof of the isolated singularity theorem (Theorem 1.3 or 1.5) is carried out by combining several steps as follows. If a weak solution $u$ has a non-removable isolated singularity at 00, then
@@
The paper is organized as follows. In Section 2, we summarize the notation and preliminary results on functional inequalities, weak solutions and capacities. Section 3 consists of local regularity estimates, namely, various Caccioppoli-type estimates, Hölder regularity and the Harnack inequality. The proofs for Theorems 1.1 and 1.3 are presented in Sections 4 and 5, respectively.
@@ 2.1. Functional inequalities
Some functional inequalities, such as the fractional Sobolev inequality and Hölder’s inequality, are presented here. We begin with fractional Sobolev inequalities. See Di Nezza–Palatucci–Valdinoci [21, Theorems 6.5 and 6.7] for their proofs.
@@
(Cozzi [17, Lemma 4.6]) Let $\tilde{s}\in(0,s)$, $\tilde{p}\in(1,p)$ and $B=B_{r}(x_{0})$. Then there exists a constant $C=C(n,s,\tilde{s},p,\tilde{p})>0$ such that
@@ 2.2. Weak solution
where $\Lambda$ is the ellipticity constant given in (1.2).
@@ 2.3. $(s,p)$-capacity
As pointed out in Björn–Björn–Kim [6, Section 5], (2.7) and (2.8) do not change the definition of capacities for compact sets.
The following estimates for the $(s,p)$-capacity of balls and annulus will be useful in the sequel.
@@
$\mathrm{cap}_{s,p}(\overline{B}_{r}\setminus B_{r/2},B_{R})\eqsim\mathrm{cap}_{s,p}(\overline{B}_{r},B_{R})\eqsim\begin{cases}\left(r^{-\frac{n-sp}{p-1}}-R^{-\frac{n-sp}{p-1}}\right)^{1-p}&\text{if }sp<n,\\ (\log(R/r))^{1-p}&\text{if }sp=n,\end{cases}$
@@
The second comparability is contained in Kim–Lee–Lee [26, Lemma 2.17], so it is enough to prove the first comparability. Since the capacity is monotone increasing with respect to the argument, it suffices to prove that
@@
Let $\Omega$ be a bounded open set and let $\varepsilon>0$. Since $\mathrm{cap}_{s,p}(E\cap\Omega,\Omega)=0$, there exists an open neighborhood $G\subset\Omega$ of $E\cap\Omega$ such that $\mathrm{cap}_{s,p}(G,\Omega)<\varepsilon$. Let $K\subset G$ be a compact set and take $\varphi\in W(K,\Omega)$ such that $[\varphi]_{W^{s,p}(\mathbb{R}^{n})}^{p}\leq\mathrm{cap}_{s,p}(K,\Omega)+\varepsilon<2\varepsilon$. Thus, the fractional Poincaré inequality (see e.g. Björn–Björn–Kim [6, (5.6)]) shows that
@@
It follows from [6, Lemma 5.5] that $sp\leq n$ if and only if $\mathrm{cap}_{s,p}(\{x_{0}\},\Omega)=0$ for every bounded open $\Omega$ containing $x_{0}$. Since $\mathrm{cap}_{s,p}(\{x_{0}\},\cdot)$ is decreasing with respect to the set inclusion, the desired result follows. ∎
@@ 3.1. Caccioppoli-type estimates
In order to state the Caccioppoli-type estimates, we fix $\beta\geq\beta_{0}>0$, $\gamma=\beta+p-1$ and $\gamma_{0}=\beta_{0}+p-1$. For $l>d>0$, we define the functions $F,G:[d,\infty)\to[0,\infty)$ by
@@
where the constants $c,C>0$ depend only on $p$.
@@
We may assume that $a>b$. Note that since $a-b\geq A-B$ and $(a-b)^{p-1}G(B)=(A-B)^{p-1}G(B)$, we have
$L$ $\coloneqq\frac{\gamma}{\beta_{0}}(a-b)^{p-1}(G(A)\eta_{1}^{p}-G(B)\eta_{2}^{p})$
$=\frac{\gamma}{\beta_{0}}(a-b)^{p-1}(G(A)-G(B))\eta_{1}^{p}+\frac{\gamma}{\beta_{0}}(a-b)^{p-1}G(B)(\eta_{1}^{p}-\eta_{2}^{p})$
$\geq\frac{\gamma}{\beta_{0}}(A-B)^{p-1}(G(A)-G(B))\eta_{1}^{p}+\frac{\gamma}{\beta_{0}}(A-B)^{p-1}G(B)(\eta_{1}^{p}-\eta_{2}^{p})$
$=\frac{\gamma}{\beta_{0}}(A-B)^{p-1}(G(A)-G(B))\eta_{2}^{p}+\frac{\gamma}{\beta_{0}}(A-B)^{p-1}G(A)(\eta_{1}^{p}-\eta_{2}^{p}).$
@@
Let us write $B_{r}=B_{r}(x_{0})$ for simplicity. We define $\varphi=G(\bar{u})(\eta\bar{\eta})^{p}\geq 0$. By Proposition 2.6, $\varphi$ can be used as a test function for (1.1) in $B_{R}\setminus E$. Thus, we have
$\mathcal{E}(u,\varphi)+\int_{B_{R}\setminus E}b(x,u(x))\varphi(x)\,\mathrm{d}x\leq 0.$
@@
Combining all the estimates above and using Young’s inequality yields
@@
Combining the estimates above yields (3.6) with the exponent $\varepsilon/(sp)$ replaced by $\varepsilon/(2\tilde{s}p)$. The desired estimate (3.3) then follows from [39, Lemma 2] and that $\gamma^{2\tilde{s}p(p-1)/\varepsilon}\leq\gamma^{p-1}+\gamma^{2sp(p-1)/\varepsilon}$ as in the previous case. ∎
@@
Note that Lemmas 3.1 and 3.3 are Moser-type Caccioppoli estimates. We also need the following De Giorgi-type Caccioppoli estimate.
@@
We basically follow the lines of proof of Proposition 8.5 in Cozzi [17] to estimate the left-hand side of (3.9), but we keep the cut-off functions $\eta$ and $\bar{\eta}$. Indeed, a slight modification of [17, (8.4)–(8.6) and (8.10)–(8.12)] shows that
@@
The rest of the proof goes the same way as before. ∎
@@
When $E=\emptyset$ in Lemma 3.4, we obtain the following standard Caccioppoli estimate. It shows that any weak subsolution of (1.1) belongs to the so-called (fractional) De Giorgi class, cf. Section 6.1 in Cozzi [17]. Note that such Caccioppoli estimates were obtained for equations similar to (1.1) in [17, Proposition 8.5], but the equation (1.1) was not covered.
@@
The proof goes as in those of Lemmas 3.4 and 3.5, with the only difference that we take $\bar{\eta}\equiv 1$. ∎